Twice Jim's age, added to Sally's age gives a total of 30. Jim's age 4 years from now will be the same as Sally's age 8 years ago. How old is each now?
Answer by checkley77(12844)

2J+S=30 OR S=30-2J
(30-2J)-8=J+4
30-2J-8=J+4
-2J-J=4+8-30
-3J=-18
J=-18/-3
J=6 JIM'S AGE NOW.
2*6+S=30
12+S=30
S=30-12
S=18 SALLY'S AGE NOW.
PROOF:
30-2*6-8=6+4
30-12-8=10
10=10